Assume that when adults with smartphones are randomly selected , 62% use them in meeting or classes. If 30 adult smartphones users are randomly selected , find the probability that exactly 22 of them use their smartphones in meetings or classes.
The probability is ?
Answer #1

Solution

Given that ,

p = 0.62

1 - p = 1 - 0.62 = 0.38

n = 30

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 22 ) = (30 C 22) * (0.62)22 * (0.38)58

Probability = 0.0689